Grade 12th PassMechanics

An engine pumps a liquid of density 'd' continuously through a pipe of area of cross-section A.If the speed with which the liquid passes through a pipe is v,then the rate of liquid flow is

a.(Adv³)/2

b.(1/2)Adv

c.(Adv²)/2

d.Adv²

Menka Malguri
14 Years agoGrade 12th Pass

1 Answer

Suresh
9 Years ago
Rate of liquid flow =(1/2*m*v^2)/tRate of flow =m/tRate of flow=dV/tRate of flow=dAL/t1/2* d*A*v *v^21/2 *dAv^3